OpenStudy (anonymous):

I believe this is correct... \[\sum_{n=1}^{\infty}\frac{x^{n}}{5^{n}n^{5}}\] \[\lim_{n\rightarrow \infty} \frac{x^{n+1}}{5^{n+1}(n+1)^{5}} \frac{5^{n}n^{5}}{x^{n}}\] \[\lim_{n\rightarrow \infty} \frac{x^{n+1}5^{n}}{x^{n}5^{n+1}} \frac{n^{5}}{(n+1)^{5}}\] \[\lim_{n\rightarrow \infty} \frac{x}{5} \frac{n^{5}}{(n+1)^{5}}\] \[\frac{1}{5}\lim_{n\rightarrow \infty} \left|x\right| \frac{n^{5}}{(n+1)^{5}}\] \[\frac{1}{5}\lim_{n\rightarrow \infty} \left|x\right| \frac{1}{(1+1/n)^{5}}\] \[\frac{1}{5}\lim_{n\rightarrow \infty} \left|x\right| \frac{1}{(1+1/n)^{5}}<1\]

OpenStudy (anonymous):

R=5

OpenStudy (anonymous):

\[\left|x\right|<5\]

OpenStudy (anonymous):

I= [-5,5]

OpenStudy (anonymous):

Radius of convergence and interval of convergence

OpenStudy (anonymous):

what do you think @agentx5

OpenStudy (anonymous):

Did you try the ratio test? This looks a lot like a combination series variant btw, a big hint to use the ratio test typically. :-D

OpenStudy (anonymous):

yes I did the ratio test

OpenStudy (anonymous):

\[\frac{a_{n+1}}{a_{n}}\]

OpenStudy (anonymous):

Ah wait I see you posted it later, ack time delay. Well since we know by the ratio test, the series converges when abs(x)<5, we also know the radius of convergence. What you need now is to find the interval of convergence. We’ll get most (if not all) of the interval in question by solving the inequality from above. Do you remember how to write that as a three-part inequality? (two inequality signs)

OpenStudy (anonymous):

I believed it was [-5,5]

OpenStudy (anonymous):

about the three part in equality ...can you show me...i'm not sure

OpenStudy (anonymous):

Right that's your radius :D But you needed the interval too right?

OpenStudy (anonymous):

Sure! one sec...

OpenStudy (anonymous):

\[-5<x<5\]

OpenStudy (anonymous):

For the less than, greater than would be...

OpenStudy (anonymous):

\[x<-5 \cup x>5\]

OpenStudy (anonymous):

so it would make more sense to write -5<x<5 instead of [-5,5]

OpenStudy (anonymous):

It does, and I would get in the practice of it because you might end up with something like this: \[ \left| x+3 \right| < 4\]

OpenStudy (anonymous):

Which is: -7<x<1

OpenStudy (anonymous):

ok what does \[x<-5 \cup x>5\] mean again

OpenStudy (anonymous):

After you do this and find this inequality, or union of inequalities, you want to determine if the power series will converge or diverge at the endpoints of this interval.

OpenStudy (anonymous):

@MathSofiya Remember the number line stuff you did way back when you were starting Algebra?

OpenStudy (anonymous):

Disjunction and conjunction?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

set of all x's less than -5 "union" set of all x's greater than +5

OpenStudy (anonymous):

Union is a fancy way of saying, "yeah both of these at the same time are true"

OpenStudy (anonymous):

i see

OpenStudy (anonymous):



OpenStudy (anonymous):

That. lol

OpenStudy (anonymous):

Versus the in-between with: 

OpenStudy (anonymous):

to determine if the power series will converge or diverge at the endpoints of this interval. do I use the initial equation? or this \[\frac{1}{5}\lim_{n\rightarrow \infty} \left|x\right| \frac{1}{(1+1/n)^{5}}\]

OpenStudy (anonymous):

You'll simply plug them at these points into the original power series and see if the series converges or diverges using any test necessary, to determine convergence. Make sense?

OpenStudy (anonymous):

yes sir

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

PS: Bump this so the others notice it and we get some additional reviewers :-)

OpenStudy (anonymous):

Did the answer check out ok?

OpenStudy (anonymous):

agent, can you help cal and on with a geometry problem once yo uare done?